A crate is lowered at a constant velocity of 2.5 ms-1 by means of a cable wrapped round a pulley. The effective diameter of the pulley 0.20m. Calculate the:

i. angular velocity of the pulley

ii. number of revolutions it makes when the crate descends 15 m.


1
Expert's answer
2021-05-30T13:22:12-0400

Solution.

"v=2.5m\/s;"

"d=0.20m;"

"i. v=\omega r\implies \omega=\dfrac{v}{r};
 \omega=\dfrac{2.5m/s}{0.10m}=25s^{-1};"

"ii. l=15m;"

"C=2\pi r;"

"C=2\sdot3.14\sdot0.1m=0.628m;"

"N=\dfrac{l}{C};"

"N=\dfrac{15m}{0.628m}=23;"


Answer: "i. 25 s^{-1};"

"ii. 23."